### v1

@@ Abstract
Context:
Aims:We explore the mass content of galaxies residing in galaxy clusters at $z=0$ in the EAGLE cosmological hydrodynamical simulation as well as the galaxies’ mass build-up through cosmic time.
Methods:We use a galaxy catalogue generated with the hbt+ algorithm, which identifies subhaloes consistently over time by tracking their dynamical evolution throughout the simulation.
Results:The satellite subhalo-to-stellar mass relation (SHSMR) is well described by a double power-law, becoming increasingly steeper with stellar mass. At stellar masses $109<\log m_{\star}/\mathrm{M}_{\odot}<$ 10, satellites have 20-25% the subhalo mass of central galaxies at fixed stellar mass. At high stellar masses, $m_{\star}>2\times 10^{11}\,\mathrm{M}_{\odot}$, the satellite SHSMR is consistent with that of centrals. The satellite SHSMR decreases steeply for satellites closer to the cluster centre, even in projection, broadly consistent with recent weak lensing measurements. The scatter in the satellite SHSMR is larger than that of central galaxies at all cluster masses and cluster-centric distances $R<R_{\mathrm{200m}}$. The SHSMR scatter decreases with stellar mass by about 12% over an order of magnitude, but this dependence can be explained by the mixing of infall times when binning by stellar mass. By splitting satellites into direct and indirect infallers (those that fell into their current host as a central galaxy or as satellite of an infalling group, respectively) we clearly show the impact of preprocessing separately on satellite galaxies’ dark and stellar mass. There is significant dark matter preprocessing; the most recent infallers into massive ($M_{\mathrm{200m}}\gtrsim 10^{14}\,\mathrm{M}_{\odot}$) clusters had already lost up to 50% of their dark matter by the time of infall, particularly if they fell in indirectly as satellites of another host. On the contrary, on average satellite galaxies are still gaining stellar mass at the time of infall and they do so for another 2 Gyr afterwards, although we see evidence of a slowing growth for indirect infallers. How much and for how long they continue to gain stellar mass depends primarily on the gas mass fraction available at infall. Overall, pre- and post-processing each have similar impacts on the satellite SHSMR. Finally, we provide a simple prescription to infer the mean mass loss experienced by satellites as a function of cluster-centric distance based on a comparison to central galaxies, convenient for observational weak lensing measurements.
Conclusions:
The history and mass content of cluster galaxies in the EAGLE simulation
Cristóbal Sifón 11    Jiaxin Han 223344
(Received XXX; accepted XXX)
@@ 1 Introduction
Over time, galaxies evolve and co-evolve as they are pulled together by gravity and merge into larger galaxies or form larger systems such as groups and clusters of galaxies. Dynamical timescales in these objects are long, so a large fraction of the galaxies that fall into them survive as satellite galaxies. As this process unfolds, tidal interactions as well as hydrodynamic processes transfer mass from the satellites to the host system, changing the composition of a galaxy depending on its environment, and the dark matter haloes hosting these galaxies become subhaloes bound to the parent structure (e.g., Gunn & Gott, 1972; White & Rees, 1978; Springel et al., 2001). The clearest observational evidence of this interaction is the rapid removal of gas from infalling galaxies and subsequent shutdown of star formation, most notably observed in fraganti in so-called “jellyfish” galaxies (Ebeling et al., 2014; Jaffé et al., 2018). Thus blue, star-forming spiral galaxies are transformed to “red and dead”, passively-evolving ellipticals. This interaction gives rise to the well-known morphology-density relation (Dressler et al., 1980) and a tight red sequence which is most prominent in massive galaxy clusters (Gladders & Yee, 2000; Koester et al., 2007).
@@
A similarly dramatic experience is expected for the dark matter halo of a galaxy driven by tidal interactions. Dark matter is more loosely bound than stars, which are heavily concentrated in the centres of galaxies, and is therefore more susceptible to tidal stripping (e.g., Chang et al., 2013; Smith et al., 2016). Thus, as a galaxy spends more time within a larger halo it progressively loses preferentially more dark matter—its total-to-stellar mass ratio decreases with time. Directly probing the dark matter content of galaxies requires gravitational lensing or spectroscopic measurements of stellar velocities within galaxies (i.e., rotation curves or velocity dispersions). Such measurements have provided tentative evidence of mass stripping using mass segregation as a proxy, i.e., taking distance from the cluster centre as a proxy for time-since-infall, and evaluating the subhalo-to-stellar mass ratio or, in some cases, the mean sizes of galaxies. Early gravitational lensing analyses suggested that galaxies are significantly more concentrated the closer they are to a cluster centre, which was taken as evidence of mass loss through stripping (e.g., Limousin et al., 2005, 2007; Natarajan et al., 2009), although it is possible that these results were at least partly driven by the modeling choices (e.g., Mira et al., 2011). More recently weak lensing analyses measuring the average total mass in subhaloes have found that the total-to-stellar mass ratio decreases with cluster-centric distance, as expected (Sifón et al., 2015, 2018; Niemiec et al., 2017; Li et al., 2016; Kumar et al., 2022; Wang et al., 2023), but a quantitative comparison between works is complicated by differing definitions of subhalo mass and density profiles (Sifón et al., 2018).
@@
The impact of tidal effects on the dark matter (sub)haloes has been assessed to great extent in dark matter only simulations, which tend to agree that dark matter subhaloes can lose as much as 90% of their mass over the timescale of a few Gyrs after entering a massive halo (e.g. Rhee et al., 2017). A significant fraction of subhaloes in $N$-body simulations are completely disrupted because of these interactions, but most of this may be attributed to numerical noise (Klypin et al., 1999a; Han et al., 2016; van den Bosch et al., 2018) or to the difficulty in accurately separating subhaloes from parent haloes (Knebe et al., 2011; Han et al., 2018), which complicates their interpretation.
@@
In comparing to observations, $N$-body simulations provide a valuable first approach but are limited when performing detailed comparisons, as highlighted in several well-known discordances (e.g., Klypin et al., 1999b; Moore et al., 1999; Boylan-Kolchin et al., 2011; Jauzac et al., 2016), which are largely alleviated when gas heating by supernovae and active galactic nuclei are incorporated into the simulations (e.g., Gao et al., 2004; Zolotov et al., 2012; Brooks et al., 2013; Bahé, 2021). Therefore, simulations that incorporate these hydrodynamical effects are required in order to inform our interpretation—be it astrophysical or cosmological—of observations in more detail. Such efforts are now possible thanks to high-resolution cosmological hydrodynamical simulations which allow predictions of galaxy observables instead of only dark matter haloes (e.g., Vogelsberger et al., 2014; Schaye et al., 2015; Pillepich et al., 2018). Indeed, recent studies have made use of hydrodynamical simulations to establish more realistic expectations for the mass content of satellite galaxies (Engler et al., 2021; Niemiec et al., 2022; O’Neil et al., 2023). Characterizing the relation between luminous and dark matter in satellite galaxies is also useful to allow the population of dark matter only simulations with galaxies with realistic properties and evolution (e.g., Lacey et al., 2016; Jiang et al., 2021; Icaza-Lizaola et al., 2023).
@@
In this paper we explore the mass content of satellite galaxies using an improved galaxy catalogue from the largest hydrodynamical simulation within the Evolution and Assembly of Galaxies and their Environments (EAGLE) project (Crain et al., 2015; Schaye et al., 2015). The catalogue was produced with the hbt+ algorithm (Han et al., 2012, 2018), which natively tracks subhaloes consistently over time. We describe the subhalo-to-stellar mass relation at $z=0$ and how it has evolved over time, paying particular attention to the mass loss experienced by infalling galaxies and the build-up of the relation between total and stellar mass through time. In this context, we discuss the
@@
This paper is organized as follows. We describe the EAGLE simulation and the galaxy and cluster samples used in Section 2. In Section 3, we look at the present-day relation between total and stellar mass and compare to available observations. In Section 4 we summarize the history of cluster galaxies and their mass content. In Section 5 we discuss the implications for the subhalo-to-stellar mass relation over time, and in Section 6 we discuss observational proxies of mass loss. We summarize our results in Section 7.
@@
Throughout, we use $\log$ to refer to base 10-logarithm. We use the notations $\langle X\mid Y\rangle$ and $\sigma(X\mid Y)$ to represent the mean and scatter in $X$ at fixed $Y$, respectively. Unless specified otherwise, masses and other properties refer to the present time, $z=0$. Lookback times are in units of Gigayears ago, expressed Gya.
@@ 2 Simulation data
We use the output of a simulation in the Evolution and Assembly of Galaxies and their Environment (EAGLE) project (Crain et al., 2015; Schaye et al., 2015). EAGLE is a suite of cosmological hydrodynamical simulations with varying box sizes, resolutions, and baryonic feedback prescriptions including supernova and AGN feedback. The simulation we use in this work is labelled RefL0100N1504 and has a box size of $(100\,\mathrm{Mpc})^{3}$, with $1504^{3}$ particles and particle masses of $1.81\times 10^{6}\,\mathrm{M}_{\odot}$ and $9.70\times 10^{6}\,\mathrm{M}_{\odot}$ for baryons and dark matter, respectively. The simulation was run assuming a flat $\Lambda$CDM cosmology with $\Omega_{\mathrm{m}}=0.307$, $\Omega_{\mathrm{b}}=0.04825$, $\sigma_{8}=0.8288$, and $h=0.6777$.
We use the subhalo catalogue produced with the upgraded Hierarchical Bound Tracing (hbt+, Han et al., 2012, 2018) post-processing of EAGLE, in which subhaloes are found consistently across time by tracing their evolution as well as their instantaneous properties. This makes hbt+ both more robust in identifying subhalos and more stable in building their evolutionary histories. In particular, hbt+ is robust against background obscuration (Han et al., 2012) that prevents most subhalo finders from identifying subhaloes near the center the host (Knebe et al., 2011; Muldrew et al., 2011). By construction, it is also immune to the flip-flop problem when linking subhaloes across time (Behroozi et al., 2015). These advantages make hbt+ an ideal choice for this work.
@@ 2.1 Mass definitions and sample
Throughout this work we speak of ‘subhaloes’ as all dark matter haloes that inhabit a larger halo, including central subhaloes (the subhaloes associated with central galaxies); and of ‘galaxies’ as the subset of subhaloes which contain at least one star particle. In practice, we limit our study to galaxies with stellar masses $m_{\star}\geq 10^{9}\,\mathrm{M}_{\odot}$. All such subhaloes have at least 1,000 particles in total, of which at least 500 stellar particles. Total masses of both subhaloes and galaxies are defined as the total mass of all bound particles (i.e., the sum of all dark matter, gas, star, and black hole particles, defined in hbt+ as Mbound), and stellar masses are similarly defined as the total mass in all bound star particles according to hbt+. Central and satellite subhaloes are hosted within a ‘host halo’, whose total mass is defined as an overdensity mass $M_{\mathrm{200m}}$ with respect to the mean density of the Universe at the corresponding redshift, contributed by the masses of all the particles within the corresponding radius, $R_{\rm 200m}$.
@@ 3 Present-day mass content of cluster galaxies in EAGLE
Figure 2: Dependence of the satellite subhalo-to-stellar mass relation on host cluster mass, $M_{\mathrm{200m}}$, as described by the colour bar, where jumps in colour mark $M_{\mathrm{200m}}$ bin edges. Shaded regions are standard errors on the means. The black dashed line with circles shows the mean SHSMR of central galaxies.
Figure 3: Total-to-stellar mass ratio (i.e., SHSMR) as a function of cluster-centric distance normalized by cluster size, $R_{\mathrm{200m}}$, binned by stellar mass. Left: 3-dimensional distances. The straight gray line shows a linear dependence with cluster-centric distance, for reference. Right: distances projected along the $z$ direction. Weak lensing measurements from Sifón et al. (2018), Kumar et al. (2022), and Wang et al. (2023) are shown with black circles, green squares, and blue triangles, respectively. Numbers in parentheses correspond to the mean stellar mass in each study, in units of $10^{10}\,\mathrm{M}_{\odot}$.
@@ 3.1 Subhalo-to-stellar mass relation
Figure 1 shows the joint distribution of subhalo and stellar masses for satellite galaxies in massive clusters in EAGLE, as well as the subhalo-to-stellar mass relation (SHSMR), i.e., $\langle m_{\mathrm{sub}}\mid m_{\star}\rangle$, for central and satellite galaxies. Following Niemiec et al. (2022) we fit the satellite and central SHSMRs separately using double power-laws of the form
@@
Best-fit parameters for centrals and satellites are listed in Table 1. The ratio of satellite-to-central SHSMR increases from $0.19\pm 0.04$ at $m_{\star}=10^{9}\,\mathrm{M}_{\odot}$ to consistent with unity at $m_{\star}>2\times 10^{11}\,\mathrm{M}_{\odot}$. As shown in Figure 1, over the range $9<\log m_{\star}/\mathrm{M}_{\odot}\lesssim 11$ the EAGLE SHSMR agrees with the SHSMR of satellite cluster galaxies in the IllustrisTNG simulation derived by Niemiec et al. (2022). At larger stellar masses the EAGLE SHSMR tends to lie above that from IllustrisTNG. This could be largely explained by the different subhalo finders used. The most massive subhaloes usually have their outer part well mixed with their host haloes, making it difficult to separate the two using only a single snapshot. In IllustrisTNG, subhaloes were identified using the (instantaneous) configuration-space subhalo finder subfind, which has difficulty determining the masses of subhaloes deep into their host haloes (e.g. Knebe et al., 2011; Onions et al., 2012; Han et al., 2018). As hbt+ uses temporal information when identifying subhalos, it is able to recover a much more complete list of member particles for these massive subhaloes. Most notably, the SHSMR for satellites at the highest mass end approaches that for central galaxies according to hbt+. This is consistent with the discussions in Han et al. (2018) that the most massive subhaloes experience weaker tidal stripping and stronger dynamical friction, so that the surviving massive ones are not much different from centrals. Nevertheless, to draw conclusions that are not sensitive to this halo-finding systematic, in the remainder of this work we study satellite galaxies over the range $119<\log m_{\star}/\mathrm{M}_{\odot}<$ 11. As we discuss next, this is also the range over which there are observational measurements to compare with.
@@
We also compare the EAGLE SHSMR with galaxy-galaxy lensing measurements of satellites galaxies in massive ($M_{\mathrm{200m}}\gtrsim 10^{14}\,\mathrm{M}_{\odot}$) clusters (Sifón et al., 2018), which is consistent with measurements at lower cluster masses (e.g., Dvornik et al., 2020), in Figure 1. The mass range probed by Sifón et al. (2018) stops right where EAGLE predicts a turn in slope so that the weak lensing measurements are not able to assess the double power law nature of the SHSMR, but over the range probed the EAGLE SHSMR agrees with the observational measurements. Figure 2 shows that the SHSMR [^1] In order to reduce the dynamical range, we represent the SHSMR as $\langle m_{\mathrm{sub}}/m_{\star}\mid m_{\star}\rangle$ hereafter. depends weakly on cluster mass. At low stellar masses, $m_{\mathrm{sub}}\lesssim 10^{10}\,\mathrm{M}_{\odot}$, there is a constant $\sim$10difference in $m_{\mathrm{sub}}/m_{\star}$ over an order of magnitude in cluster mass. At higher stellar masses. $10^{10}\lesssim m_{\star}/\mathrm{M}_{\odot}\lesssim 10^{11}$, galaxies have approximately twice as much total mass at fixed stellar mass in $M_{\mathrm{200m}}\sim 10^{13}\,\mathrm{M}_{\odot}$ than in $M_{\mathrm{200m}}\sim 10^{14}\,\mathrm{M}_{\odot}$ clusters, although this trend seems to stall for larger cluster masses. That this difference manifests itself only for more massive satellites can be generally ascribed to their shorter dynamical timescales. The trend with cluster mass, on the other hand, can be due to two effects, noting that—as discussed later—dark matter is more susceptible to mass loss than stellar mass. Firstly, tidal stripping is expected to be stronger in more massive clusters. Secondly, galaxies in more massive clusters are expected to have been more strongly preprocessed, such that they enter massive clusters already with a reduced dark matter fraction (Bahé et al., 2019). We return to this in Section 5.

### v2

@@ Abstract
Aims. We explore the mass content of galaxies residing in galaxy clusters at $z=0$ in the EAGLE cosmological hydrodynamical simulation. We also explore the galaxies’ mass build-up through cosmic time.
Methods. We used a galaxy catalogue generated with the hbt+ algorithm, which identifies subhaloes consistently over time by tracking their dynamical evolution throughout the simulation.
Results. The satellite subhalo-to-stellar mass relation (SHSMR) is well described by a double power law, becoming increasingly steeper with stellar mass. At stellar masses $109<\log m_{\star}/\mathrm{M}_{\odot}<$ 10, satellites have 20-25% the subhalo mass of central galaxies at fixed stellar mass. At high stellar masses, $m_{\star}>2\times 10^{11}\,\mathrm{M}_{\odot}$, the satellite SHSMR is consistent with that of centrals. The satellite SHSMR decreases steeply for satellites closer to the cluster centre, even in projection, broadly consistent with recent weak lensing measurements. The scatter in the satellite SHSMR is larger than that of central galaxies at all cluster masses and cluster-centric distances $R<R_{\mathrm{200m}}$. The SHSMR scatter decreases with stellar mass by about 12% over an order of magnitude, but this dependence can be explained by the mixing of infall times when binning by stellar mass. By splitting satellites into direct and indirect infallers (those that fell into their current host as a central galaxy or as the satellite of an infalling group, respectively) we clearly show the impact of pre-processing separately on satellite galaxies’ dark and stellar mass. There is significant dark matter pre-processing; the most recent infallers into massive clusters ($M_{\mathrm{200m}}\gtrsim 10^{14}\,\mathrm{M}_{\odot}$) had already lost up to 50% of their dark matter by the time of infall, particularly if they fell in indirectly as satellites of another host. On the contrary, on average, satellite galaxies are still gaining stellar mass at the time of infall and they do so for another 2 Gyr afterwards, although we see evidence of a slowing growth for indirect infallers. How much and for how long they continue to gain stellar mass depends primarily on the gas mass fraction available at infall. Overall, pre- and post-processing have similar impacts on the satellite SHSMR. Finally, we provide a simple prescription to infer the mean mass loss experienced by satellites as a function of cluster-centric distance based on a comparison to central galaxies, convenient for observational weak lensing measurements.
@@ 1 Introduction
Over time, galaxies evolve and co-evolve as they are pulled together by gravity and merge into larger galaxies or form larger systems, such as groups and clusters of galaxies. Dynamical timescales in these objects are long, so a large fraction of the galaxies that fall into them survive as satellite galaxies. As this process unfolds, tidal interactions as well as hydrodynamic processes transfer mass from the satellites to the host system, changing the composition of a galaxy depending on its environment, and the dark matter haloes hosting these galaxies become subhaloes bound to the parent structure (e.g. Gunn & Gott, 1972; White & Rees, 1978; Springel et al., 2001). Both gravitational and hydrodynamical interactions then unfold, transforming blue star-forming spiral galaxies into ‘red and dead’ passively evolving ellipticals, giving rise to the well-known morphology-density relation (Dressler et al., 1980) and a tight red sequence that is most prominent in massive galaxy clusters (Gladders & Yee, 2000; Koester et al., 2007). The clearest observational evidence of these interactions is found in ‘jellyfish’ galaxies, in which radio observations reveal gas being removed from infalling galaxies, sometimes accompanied by trails of star-forming regions observable in optical wavelengths (Ebeling et al., 2014; Jaffé et al., 2018).
@@
A similarly dramatic experience is expected for the dark matter halo of a galaxy driven by tidal interactions. Dark matter is more loosely bound than stars are (and stars are heavily concentrated in the centres of galaxies), and is therefore more susceptible to tidal stripping (e.g. Chang et al., 2013; Smith et al., 2016). Thus, as a galaxy spends more time within a larger halo it progressively and preferentially loses more dark matter; its total-to-stellar mass ratio decreases with time. Directly probing the dark matter content of galaxies requires gravitational lensing or spectroscopic measurements of stellar velocities within galaxies (i.e. rotation curves or velocity dispersions). Such measurements have provided tentative evidence of mass stripping by evaluating changes in galaxies’ mass content or, in some cases, their mean sizes, with distance from the cluster centre as a proxy for time-since-infall. Early gravitational lensing analyses suggested that galaxies are significantly more concentrated the closer they are to a cluster centre, which was taken as evidence of mass loss through stripping (e.g. Limousin et al., 2005, 2007; Natarajan et al., 2009), although it is possible that these results were at least partly driven by the modelling choices (e.g. Mira et al., 2011). More recently weak lensing analyses measuring the average total mass in subhaloes have found that the total-to-stellar mass ratio decreases with cluster-centric distance, as expected (Sifón et al., 2015, 2018; Niemiec et al., 2017; Li et al., 2016; Kumar et al., 2022; Wang et al., 2023), but a quantitative comparison between works is complicated by differing definitions of subhalo mass and density profiles (Sifón et al., 2018). Because we base our comparison to observational work on the aforementioned lensing studies, and because tidal stripping is expected to directly shape the subhalo-to-stellar mass ratio (Han et al., 2016), in this paper we use the term ‘mass segregation’ to mean changes in the subhalo-to-stellar mass relation as done by those authors, although we note that this term is sometimes used to mean changes in the mean stellar or subhalo masses alone (e.g. Joshi et al., 2017; Kim2020).
@@
The impact of tidal effects on the dark matter (sub)haloes has been assessed to great extent in dark matter-only simulations, which tend to agree that dark matter subhaloes can lose as much as 90% of their mass over timescale of a few Gyrs after entering a massive halo (e.g. Rhee et al., 2017). A significant fraction of subhaloes in $N$-body simulations are completely disrupted because of these interactions, but most of this may be attributed to numerical noise (Klypin et al., 1999a; Han et al., 2016; van den Bosch et al., 2018) or to the difficulty in accurately separating subhaloes from parent haloes (Knebe et al., 2011; Han et al., 2018), which complicates their interpretation.
@@
When comparing to observations, $N$-body simulations provide a valuable first approach, but are limited when performing detailed comparisons, as highlighted in several well-known discordances (e.g. Klypin et al., 1999b; Moore et al., 1999; Boylan-Kolchin et al., 2011; Jauzac et al., 2016), which are largely alleviated when gas heating by supernovae and active galactic nuclei are incorporated into the simulations (e.g. Gao et al., 2004; Zolotov et al., 2012; Brooks et al., 2013; Bahé, 2021). Therefore, simulations that incorporate these hydrodynamical effects are required in order to inform our interpretation (be it astrophysical or cosmological) of observations in more detail. Such efforts are now possible thanks to high-resolution cosmological hydrodynamical simulations, which allow predictions of galaxy observables instead of only dark matter haloes (e.g. Vogelsberger et al., 2014; Schaye et al., 2015; Pillepich et al., 2018). Recent studies have made use of hydrodynamical simulations to establish more realistic expectations for the mass content of satellite galaxies (Engler et al., 2021; Niemiec et al., 2022; O’Neil et al., 2023). Characterizing the relation between luminous and dark matter in satellite galaxies is also useful to allow the population of dark matter-only simulations with galaxies with realistic properties and evolution (e.g. Lacey et al., 2016; Jiang et al., 2021; Icaza-Lizaola et al., 2023).
@@
In this paper we explore the mass content of satellite galaxies using an improved galaxy catalogue from the largest hydrodynamical simulation within the Evolution and Assembly of Galaxies and their Environments (EAGLE) project (Crain et al., 2015; Schaye et al., 2015). The catalogue was produced with the hbt+ algorithm (Han et al., 2012, 2018), which natively tracks subhaloes consistently over time. We describe the subhalo-to-stellar mass relation at $z=0$ and how it has evolved over time, paying particular attention to the mass loss experienced by infalling galaxies and the build-up of the relation between total and stellar mass through time.
@@
This paper is organized as follows. We describe the EAGLE simulation and the galaxy and cluster samples used in Section 2. In Section 3 we look at the present-day relation between total and stellar mass and compare to available observations. In Section 4 we summarize the history of cluster galaxies and their mass content. In Section 5 we discuss the implications for the subhalo-to-stellar mass relation over time, and in Section 6 we discuss observational proxies of mass loss. We summarize our results in Section 7.
@@
Throughout, we use $\log$ to refer to base 10-logarithm. We use the notations $\langle X\mid Y\rangle$ and $\sigma(X\mid Y)$ to represent the mean and scatter in $X$ at fixed $Y$, respectively. Unless specified otherwise, masses and other properties refer to the present time, $z=0$. Lookback times are in units of gigayears ago, expressed as Gya.
@@ 2 Simulation data
We used the output of a simulation in the Evolution and Assembly of Galaxies and their Environment (EAGLE) project (Crain et al., 2015; Schaye et al., 2015). EAGLE is a suite of cosmological hydrodynamical simulations with varying box sizes, resolutions, and baryonic feedback prescriptions including supernova and AGN feedback. The simulation we use in this work is labelled RefL0100N1504 and has a box size of $(100\,\mathrm{Mpc})^{3}$, with $1504^{3}$ particles and particle masses of $1.81\times 10^{6}\,\mathrm{M}_{\odot}$ and $9.70\times 10^{6}\,\mathrm{M}_{\odot}$ for baryons and dark matter, respectively. The simulation was run assuming a flat $\Lambda$CDM cosmology with $\Omega_{\mathrm{m}}=0.307$, $\Omega_{\mathrm{b}}=0.04825$, $\sigma_{8}=0.8288$, and $h=0.6777$.
We used the subhalo catalogue produced with the upgraded Hierarchical Bound Tracing (hbt+, Han et al., 2012, 2018) post-processing of EAGLE, in which subhaloes are found consistently across time by tracing their evolution as well as their instantaneous properties. This makes hbt+ both more robust in identifying subhaloes and more stable in building their evolutionary histories. In particular, hbt+ is robust against background obscuration (Han et al., 2012), which prevents most subhalo finders from identifying subhaloes near the centre of the host (Knebe et al., 2011; Muldrew et al., 2011). By construction, it is also immune to the flip-flop problem when linking subhaloes across time (Behroozi et al., 2015). These advantages make hbt+ an ideal choice for this work.
@@ 2.1 Mass definitions and sample
Throughout this work we speak of ‘subhaloes’ as all dark matter haloes that inhabit a larger halo, including central subhaloes (the subhaloes associated with central galaxies); and of ‘galaxies’ as the subset of subhaloes which contain at least one star particle. In practice, we limit our study to galaxies with stellar masses $m_{\star}\geq 10^{9}\,\mathrm{M}_{\odot}$. All such subhaloes have at least 1,000 particles in total, at least 500 of which are stellar particles. Total masses of both subhaloes and galaxies are defined as the total mass of all bound particles (i.e. the sum of all dark matter, gas, star, and black hole particles, defined in hbt+ as Mbound), and stellar masses are similarly defined as the total mass in all bound star particles according to hbt+. Central and satellite subhaloes are hosted within a ‘host halo’, whose total mass is defined as an overdensity mass $M_{\mathrm{200m}}$ with respect to the mean density of the Universe at the corresponding redshift, contributed by the masses of all the particles within the corresponding radius, $R_{\rm 200m}$.
@@
In some cases it is useful to compare the properties of satellites to those of centrals of the same stellar mass. With the stellar mass cut of $119<\log m_{\star}/\mathrm{M}_{\odot}<$ 11, there are in total 7,205 centrals in the simulation box at $z=0$. These centrals reside mostly in lower mass haloes than clusters, with masses typically in the range $12.511<\log M_{\mathrm{200m}}<$ 12.5.
@@ 3 Present-day mass content of cluster galaxies in EAGLE
Figure 2: Dependence of the satellite subhalo-to-stellar mass relation on host cluster mass, $M_{\mathrm{200m}}$, as described by the colour bar, where jumps in colour mark $M_{\mathrm{200m}}$ bin edges. The shaded regions are standard errors on the means. The black dashed line with circles shows the mean SHSMR of central galaxies.
Refer to captionRefer to caption
Figure 3: Total-to-stellar mass ratio (SHSMR) as a function of cluster-centric distance normalized by cluster size, $R_{\mathrm{200m}}$, binned by stellar mass. Horizontal lines in the top left show the mean SHSMR for centrals in each stellar mass bin. Left: 3D distances. The straight grey line shows a linear dependence with cluster-centric distance, for reference. Right: Distances projected along the $z$ direction. Weak lensing measurements from Sifón et al. (2018), Kumar et al. (2022), and Wang et al. (2023) are shown with black circles, green squares, and blue triangles, respectively. Numbers in parentheses correspond to the mean stellar mass in each study, in units of $10^{10}\,\mathrm{M}_{\odot}$.
@@ 3.1 Subhalo-to-stellar mass relation
Figure 1 shows the joint distribution of subhalo and stellar masses for satellite galaxies in massive clusters in EAGLE, as well as the subhalo-to-stellar mass relation (SHSMR), $\langle m_{\mathrm{sub}}\mid m_{\star}\rangle$, for central and satellite galaxies. Following Niemiec et al. (2022) we fit the satellite and central SHSMRs separately using double power laws of the form
@@
Best-fit parameters for centrals and satellites are listed in Table 1. The ratio of satellite-to-central SHSMR increases from $0.19\pm 0.04$ at $m_{\star}=10^{9}\,\mathrm{M}_{\odot}$ to consistent with unity at $m_{\star}>2\times 10^{11}\,\mathrm{M}_{\odot}$. As shown in Figure 1, over the range $9<\log m_{\star}/\mathrm{M}_{\odot}\lesssim 11$ the EAGLE SHSMR agrees with the SHSMR of satellite cluster galaxies in the IllustrisTNG simulation derived by Niemiec et al. (2022). At larger stellar masses the EAGLE SHSMR tends to lie above that from IllustrisTNG. This could be largely explained by the different subhalo finders used. The most massive subhaloes usually have their outer part well mixed with their host haloes, making it difficult to separate the two using only a single snapshot. In IllustrisTNG, subhaloes were identified using the (instantaneous) configuration-space subhalo finder subfind, which has difficulty determining the masses of subhaloes deep into their host haloes (e.g. Knebe et al., 2011; Onions et al., 2012; Han et al., 2018). As hbt+ uses temporal information when identifying subhaloes, it is able to recover a much more complete list of member particles for these massive subhaloes. Most notably, the SHSMR for satellites at the highest mass end approaches that for central galaxies according to hbt+. This is consistent with the discussions in Han et al. (2018) that the most massive subhaloes experience weaker tidal stripping and stronger dynamical friction, so that the surviving massive ones are not much different from centrals. Nevertheless, to draw conclusions that are not sensitive to this halo-finding systematic, in the remainder of this work we study satellite galaxies over the range $119<\log m_{\star}/\mathrm{M}_{\odot}<$ 11. As we discuss next, this is also the range over which there are observational measurements to compare with.
@@
We also compare the EAGLE SHSMR with galaxy-galaxy lensing measurements of satellites galaxies in massive ($M_{\mathrm{200m}}\gtrsim 10^{14}\,\mathrm{M}_{\odot}$) clusters (Sifón et al., 2018), which is consistent with measurements at lower cluster masses (e.g. Dvornik et al., 2020), in Figure 1. The stellar mass range probed by Sifón et al. (2018) stops right where EAGLE predicts a turn in slope so that the weak lensing measurements are not able to assess the double power-law nature of the SHSMR, but over the range probed the EAGLE SHSMR agrees with the observational measurements. Figure 2 shows that the SHSMR [^1] In order to reduce the dynamical range, we represent the SHSMR as $\langle m_{\mathrm{sub}}/m_{\star}\mid m_{\star}\rangle$ hereafter. depends weakly on cluster mass. At low stellar masses, $m_{\mathrm{sub}}\lesssim 10^{10}\,\mathrm{M}_{\odot}$, there is a constant $\sim$10difference in $m_{\mathrm{sub}}/m_{\star}$ over an order of magnitude in cluster mass. At higher stellar masses, $10^{10}\lesssim m_{\star}/\mathrm{M}_{\odot}\lesssim 10^{11}$, galaxies have approximately twice as much total mass at fixed stellar mass in $M_{\mathrm{200m}}\sim 10^{13}\,\mathrm{M}_{\odot}$ than in $M_{\mathrm{200m}}\sim 10^{14}\,\mathrm{M}_{\odot}$ clusters, although this trend seems to stall for larger cluster masses. That this difference manifests itself only for more massive satellites can be generally ascribed to their shorter dynamical timescales. The trend with cluster mass, on the other hand, can be due to two effects, noting that—as discussed later—dark matter is more susceptible to mass loss than stellar mass. Firstly, tidal stripping is expected to be stronger in more massive clusters. Secondly, galaxies in more massive clusters are expected to have been more strongly pre-processed, such that they enter massive clusters already with a reduced dark matter fraction (Bahé et al., 2019). We return to this in Section 5.